# Asymptotic Behavior of Rupture Solutions for the Elliptic MEMS Equation with Hénon-Type Term

### v1

@@ 4 Existence of Solutions
admits a bounded inverse. We begin with the Dirichlet boundary-value problem
@@
Proof. We give the argument for the case $\delta^{(k_{0}-1)}<\alpha<\delta^{(k_{0})}$. Let $g=0$ and $\varphi=0$, suppose $v\in C_{\mu}^{2}\!\left((-\infty,t_{0}]\times S^{N-1}\right)$ is a solution of the problem. For each $k\geq 0$, define
@@
Applying the Green identity, we obtain
@@
For $k\geq k_{0}$, the coefficient
$\lambda_{k}-(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)>0.$
Thus $v_{k}(t)\equiv 0$. Therefore $v\equiv 0$.
@@
Proof. Fix an arbitrary $t\leq t_{0}$ and consider two cases.
(i) $t<t_{0}-2$. By the interior Schauder estimate,
@@
To estimate the Hölder seminorm of $g$ on $t-2,t+2]\times S^{N-1}$, take arbitrary $(t_{1},\theta_{1}),(t_{2},\theta_{2})\in[t-2,t+2]\times S^{N-1}$ with $(t_{1},\theta_{1})\neq(t_{2},\theta_{2})$. Then we split into two cases: $|t_{1}-t_{2}|\leq 2$ and $|t_{1}-t_{2}|>2$. We present the proof of the first case.
If $|t_{1}-t_{2}|\leq 2$, then there exists some $t^{\prime}\in[t-1,t+1]$ such that $t_{1},t_{2}\in[t^{\prime}-1,t^{\prime}+1]$. Hence,
@@
$\quad\leq C\left[\|v\|_{L^{\infty}\left([t-2,t+2]\times S^{N-1}\right)}+\|g\|_{L^{\infty}\left([t-2,t+2]\times S^{N-1}\right)}+\sup_{t^{\prime}\in[t-1,t+1]}[g]_{C^{\alpha}\left(\left[t^{\prime}-1,t^{\prime}+1\right]\times S^{N-1}\right)}\right]$
@@
(ii) $t_{0}-2\leq t\leq t_{0}$. By the boundary Schauder estimate,
$\sum_{j=0}^{2}\sup_{S^{N-1}}\left|\nabla^{j}v(t,\cdot)\right|+\left[\nabla^{2}v\right]_{C^{\alpha}([t_{0}-3,t_{0}]\times S^{N-1})}$
$\qquad\leq C\left[\|v\|_{L^{\infty}([t_{0}-4,t_{0}]\times S^{N-1})}+\|g\|_{L^{\infty}([t_{0}-4,t_{0}]\times S^{N-1})}+[g]_{C^{\alpha}([t_{0}-4,t_{0}]\times S^{N-1})}+\|\varphi\|_{C^{2,\alpha}(S^{N-1})}\right].$
Using an argument similar to the one above, we obtain
$\sum_{j=0}^{2}\sup_{t\in[t_{0}-2,t_{0}]}\sup_{S^{N-1}}e^{-\mu t}\left|\nabla^{j}v(t,\cdot)\right|+\sup_{t\in[t_{0}-2,t_{0}-1]}e^{-\mu t}\left[\nabla^{2}v\right]_{C^{\alpha}([t-1,t+1]\times S^{N-1})}$
$\qquad\leq C\left[\|v\|_{C_{\mu}^{0}((-\infty,t_{0}]\times S^{N-1})}+\|g\|_{C_{\mu}^{0,\alpha}((-\infty,t_{0}]\times S^{N-1})}+e^{-\mu t_{0}}\|\varphi\|_{C^{2,\alpha}(S^{N-1})}\right].$
Combining the two cases above yields the desired estimate
@@
and for every $t\in[T,t_{0}]$ and every $k=0,1,2,\ldots,K$
@@
Proof. Assume the conclusion is false. Then there exist sequences $\{T_{i}\}$, $\{t_{i}\}$, $\{v_{i}\}$ and $\{g_{i}\}$ with $t_{i}\leq 0$ and $T_{i}-t_{i}\leq-4$, such that
@@
$|\nabla\widetilde{v}_{i}|\leq Ce^{\mu(t_{i}^{*}-T_{i})}\qquad\text{on }(T_{i}-t_{i}^{*},\,T_{i}-t_{i}^{*}+1)\times S^{N-1}.$
@@
Also $\widetilde{g}_{i}$ converges uniformly to 0 on every compact subset of $(\tau_{-},\tau_{+})\times S^{N-1}$. Therefore, $\widehat{v}\neq 0$, and
$|e^{-\mu t}\widehat{v}(t,\theta)|\leq 1,$ (59)
for all $(t,\theta)\in(\tau_{-},\tau_{+})\times S^{N-1}$ .Also we have
$\mathcal{L}\widehat{v}=0.$
on $(\tau_{-},\tau_{+})\times S^{N-1}$
Furthermore,
$\lim\widehat{v}(t,\theta)=0.$ (60)
@@
$|e^{-\mu t}\widehat{v}_{k}(t)|\leq C.$
for all $t\in(\tau_{-},\tau_{+})$.
If $\tau_{+}=+\infty$, then necessarily $c_{k}^{1}=0$, thus $\widehat{v}_{k}(t)=c_{k}^{2}e^{\sigma_{2}^{(k)}t}$ which decays exponentially as $t\to+\infty$. If $\tau_{+}$ is finite,then $\lim_{t\to\tau_{+}}\widehat{v}_{k}(t)=0$ by (60),
Similarly, if $\tau_{-}=$-$\infty$, then $c_{k}^{2}=0$, hence $\widehat{v}_{k}(t)=c_{k}^{1}e^{\sigma_{1}^{(k)}t}$ which decays exponentially as $t\to$-$\infty$. If $\tau_{-}$ is finite, then $\lim_{t\to\tau_{-}}\widehat{v}_{k}(t)=0$ by (60),
@@
$\lambda_{k}-(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)>0,$
@@
$\widehat{v}_{k}\equiv 0\qquad\text{for all }k=0,1,\ldots,K.$
Combining the above, we conclude that $\widehat{v}_{k}\equiv 0$ for all $k\geq 0$, therefore $\widehat{v}\equiv 0$, contradicting our earlier conclusion that $\widehat{v}\neq 0$. Lemma 4.3 is proved.
@@
Proof. For each $k=0,1,\ldots,K$, define
@@
We claim that (61) admits a solution $v_{k}\in C_{\mu}^{2,\alpha}((-\infty,t_{0}])$ satisfying
[…]
where the constant $C$ depends only on $N,\alpha,\mu$ and not on $t_{0}$. set
@@
$v_{k}(t)$ $=B_{k}^{1}\cos\bigl(I(\sigma_{1}^{(k)}t)\bigr)\int_{-\infty}^{t}e^{\sigma_{1}^{(k)}(t-s)}g_{k}(s)\sin\bigl(I(\sigma_{1}^{(k)}s)\bigr)\,ds$ (62)
$\quad-B_{k}^{2}\cos\bigl(I(\sigma_{1}^{(k)}t)\bigr)\int_{-\infty}^{t}e^{\sigma_{2}^{(k)}(t-s)}g_{k}(s)\sin\bigl(I(\sigma_{1}^{(k)}s)\bigr)\,ds,$
$\left|B_{k}^{1}\right|=\left|\frac{1}{\sigma_{2}^{(k)}-\sigma_{1}^{(k)}}\right|.$We now write down the argument for the case of (61).
@@
Combining the above estimates yields (62).
@@
$\|v\|_{C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})}$ $\leq C\sum_{k=0}^{K}\|v_{k}\|_{C_{\mu}^{2,\alpha}((-\infty,t_{0}])}$
$\leq C\sum_{k=0}^{K}\|g_{k}\|_{C_{\mu}^{0,\alpha}((-\infty,t_{0}])}$
$\leq C\|g\|_{C_{\mu}^{0,\alpha}((-\infty,t_{0}]\times S^{N-1})}.$
Thus $v$ is the desired solution. Moreover, it’s clear that such a solution is unique under the additional condition that $v(t,\cdot)\in\operatorname{span}\{Q_{0},Q_{1},\ldots,Q_{K}\}\text{for every }t\leq t_{0}.$
@@
Proof. Fix any $T\leq t_{0}-4$. We first show that there exists $v_{T}\in C^{2,\alpha}([T,t_{0}]\times S^{N-1})$ such that
$\begin{cases}\mathcal{L}v_{T}=g&\text{in }(T,t_{0})\times S^{N-1},\\[2.84526pt] v_{T}=0&\text{on }(\{T\}\cup\{t_{0}\})\times S^{N-1}.\end{cases}$ (64)
@@
$\begin{cases}\frac{\partial}{\partial t}\left(e^{\hat{A}t}\frac{\partial v_{T}}{\partial t}\right)+e^{\hat{A}t}\Delta_{\theta}v_{T}+(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)e^{\hat{A}t}v_{T}=e^{\hat{A}t}g&\text{ in }\left(T,t_{0}\right)\times S^{N-1},\\ v_{T}=0&\text{ on }\left(\{T\}\cup\{t_{0}\}\right)\times S^{N-1},\end{cases}$
Here $\hat{A}=N-2+\frac{2\alpha+4}{p+1}$. Consider the energy functional
$\mathcal{G}_{T}(v)=\int_{T}^{t_{0}}\int_{S^{N-1}}\left[e^{\hat{A}t}(\partial_{t}v)^{2}+e^{\hat{A}t}|\nabla_{\theta}v|^{2}-(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)e^{\hat{A}t}v^{2}+2e^{\hat{A}t}gv\right]dt\,d\theta.$
@@
$\Gamma=\left\{\phi\in H^{1}(S^{N-1}):\int_{S^{N-1}}\phi(\theta)\,Q_{k}(\theta)\,d\theta=0,\quad k=0,1,2,\ldots,K,\ \sigma_{1}^{(k)}<\mu\right\}.$
Since
$\int_{S^{N-1}}|\nabla_{\theta}\phi|^{2}\,d\theta\geq K(N+K-2)\int_{S^{N-1}}\phi^{2}\,d\theta,$
it follows that for any $v\in H_{0}^{1}((T,t_{0})\times S^{N-1})$ satisfying $v(t,\cdot)\in\Gamma$ ( $t\in(T,t_{0})$), we have
$\mathcal{G}_{T}(v)\geq\int_{T}^{t_{0}}\int_{S^{N-1}}e^{\hat{A}t}(\partial_{t}v)^{2}+\left(K(N+K-2)-(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)\right)e^{\hat{A}t}v^{2}+2e^{\hat{A}t}gv~~dtd\theta.$
Since $\sigma_{1}^{(K)}>0$, we have
$K(N+K-2)-(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)>0.$
Hence, the functional $\mathcal{G}_{T}$ is coercive and weakly lower semicontinuous. Therefore we may find a minimizer $v_{T}$ of $\mathcal{G}_{T}$ in the space
$\left\{v\in H_{0}^{1}((T,t_{0})\times S^{N-1}):v(t,\cdot)\in\Gamma\text{ for every }t\in(T,t_{0})\right\}.$
Since $g(t,\cdot)\in\Gamma$ for all $t\in(T,t_{0})$, it follows that $v_{T}$ is a solution of (64), and $v_{T}(t,\cdot)\in\Gamma$ for all $t\in(T,t_{0})$.
@@
$\sup_{(t,\theta)\in[T,t_{0}]\times S^{N-1}}e^{-\mu t}\,|v_{T}(t,\theta)|\leq C\sup_{(t,\theta)\in[T,t_{0}]\times S^{N-1}}e^{-\mu t}\,|g(t,\theta)|,$
@@
For any fixed $T_{0}<t_{0}$, consider the region $t_{0}+T_{0},\,t_{0}]\times S^{N-1}\\ \subset[t_{0}+T_{0}-1,\,t_{0}]\times S^{N-1}$. By the interior and boundary Schauder estimates, together with the fact that $v_{T}(t_{0},\theta)=0$, we may extract a subsequence $v_{T}$ converges to a $C^{2,\alpha}$ solution $v$ of (55) on $t_{0}+T_{0},\,t_{0}]\times S^{N-1}$ (via the Arzelà–Ascoli theorem), with $v=0$ on $\{t_{0}\}\times S^{N-1}$ as $T\to$-$\infty$. Via a diagonalization process, we conclude that $v_{T}$ converges to a $C^{2,\alpha}$ solution $v$ of (55)on $(-\infty,t_{0}]\times S^{N-1}$ , with $v=0$ on $\{t_{0}\}\times S^{N-1}$.
@@
Substituting (65) into (57) (with $\varphi=0$) and then completes the proof.
@@
Proof. Let $K\geq 0$ be the largest integer such that $\sigma_{1}^{(K)}<\mu$. For $k=0,1,\ldots,K$, define
@@
Let $v_{1}\in C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})$ be the unique solution (by Lemma 4.4) of
@@
Next, by Lemma 3.6, choose $v_{2}\in C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})$ to be the unique solution of
@@
Then $v=v_{1}+v_{2}$ is the desired solution. Note that
$v(t_{0},\theta)=v_{1}(t_{0},\theta)=\sum_{k=0}^{K}v_{k}(t_{0})\,Q_{k}(\theta),$
where each $v_{k}(t)$ ($k=0,1,\ldots,K$) is obtained from Lemma 4.4.
Remark 4.7.
We denote by $\mathcal{L}^{-1}$ the correspondence $g\mapsto v$ in Theorem 4.6. Then
@@
It’s equivalent to prove that the equation $H(z)=0$ admits a solution.
@@
$\big|\mathcal{N}(\hat{z})\big|+\left|\frac{d}{dt}\mathcal{N}(\hat{z})\right|\leq Ce^{\mu t}\quad\forall\,t\in(-\infty,0].$
@@
Proof. We only prove case (2), since the radial case (1) is similar and easier. For any $\phi\in C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})$, we have
$\mathcal{N}(\widehat{z}+\phi)=\mathcal{N}(\widehat{z})+\mathcal{L}\phi-P(\phi),$
@@
Thus $\mathcal{N}(\widehat{z}+\phi)=0$ is equivalent to
$\mathcal{L}\phi=-\big[\mathcal{N}(\widehat{z})-P(\phi)\big].$ (67)
By Theorem 3.7 and Remark 3.8, we may rewrite (66) in the form
$\phi=\mathcal{L}^{-1}\big(-\mathcal{N}(\widehat{z})+P(\phi)\big).$
@@
$\mathcal{T}(\phi)=\mathcal{L}^{-1}\big(-\mathcal{N}(\widehat{z})+P(\phi)\big).$
@@
$\|\mathcal{N}(\widehat{z})\|_{C_{\mu}^{1}((-\infty,t_{0}]\times S^{N-1})}\leq C_{1}.$
@@
(2)Nonradial case. Assume that the corresponding index sets $\mathcal{I}_{\rho}$ and $\mathcal{I}_{\widetilde{\rho}}$ given in (37) and (38) (with multiple number ignored) in the proof of Theorem 2.2 respectively. $\mu\notin\mathcal{I}_{\rho}\cup\mathcal{I}_{\widetilde{\rho}}~\text{and}~\mu>\rho_{j+4}>\frac{2p-\alpha}{p+1}$ in cases (31)(32)(34)(35); $\mu>\rho_{1}$in cases (29)(30)(33)(36). Let $\varphi$ be a solution of the equation $\mathcal{L}(\varphi)=0$ on $\mathbb{R}\times\mathbb{S}^{N-1}$, and suppose that $\varphi(t,\theta)\to 0~~\text{as}~t\to-\infty~\text{uniformly}~\text{for}~\theta\in{S^{N-1}}$. Then there exists $t_{0}<0$ with $|t_{0}|$ sufficiently large, such that there exists a smooth function $\widetilde{\varphi}$ on $(-\infty,t_{0}]\times\mathbb{S}^{N-1}$ satisfying
@@
Proof. We prove case (2); the proof of case (1) is similar and eaiser.

### v3

@@ 4 Existence of Solutions
admits a bounded inverse. We begin with the problem
@@
Proof.
We give the argument for the case $\delta^{(k_{0}-1)}<\alpha<\delta^{(k_{0})}$. Let $g=0$ and $\varphi=0$, suppose $v\in C_{\mu}^{2}\!\left((-\infty,t_{0}]\times S^{N-1}\right)$ is a solution of the problem. For each $k\geq 0$, define
@@
Obtain by Green identity that
@@
It’s clear that the coefficient $\lambda_{k}-(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)>0$ when $k\geq k_{0}$, thus $v_{k}(t)\equiv 0$, $v\equiv 0$. ∎
@@
Proof.
Fix a $t\leq t_{0}$ and consider two cases, (i) $t<t_{0}-2$, (ii) $t_{0}-2\leq t\leq t_{0}$. They are proved by the interior and boundary Schauder estimate respectively.
For case (i), By the interior Schauder estimate,
@@
To estimate the Hölder seminorm of $g$ on $t-2,t+2]\times S^{N-1}$, take arbitrary $(t_{1},\theta_{1}),(t_{2},\theta_{2})\in[t-2,t+2]\times S^{N-1}$ with $(t_{1},\theta_{1})\neq(t_{2},\theta_{2})$. Then we split into two subcases: $|t_{1}-t_{2}|\leq 2$ and $|t_{1}-t_{2}|>2$. We present the proof of the first subcase and the other follows directly.
For $|t_{1}-t_{2}|\leq 2$, there exists some $t^{\prime}\in[t-1,t+1]$ such that $t_{1},t_{2}\in[t^{\prime}-1,t^{\prime}+1]$. Hence,
@@
$\quad\leq C\left[\|v\|_{L^{\infty}\left([t-2,t+2]\times S^{N-1}\right)}+\|g\|_{L^{\infty}\left([t-2,t+2]\times S^{N-1}\right)}+\sup_{t^{\prime}\in[t-1,t+1]}[g]_{C^{\alpha}\left(\left[t^{\prime}-1,t^{\prime}+1\right]\times S^{N-1}\right)}\right].$
@@
For case (ii), it’s similar with case (i) as we know by the boundary Schauder estimate that
$\sum_{j=0}^{2}\sup_{S^{N-1}}\left|\nabla^{j}v(t,\cdot)\right|+\left[\nabla^{2}v\right]_{C^{\alpha}\left(\left[t_{0}-3,t_{0}\right]\times S^{N-1}\right)}$
$\quad\leq C\left[\|v\|_{L^{\infty}\left(\left[t_{0}-4,t_{0}\right]\times S^{N-1}\right)}+\|g\|_{L^{\infty}\left(\left[t_{0}-4,t_{0}\right]\times S^{N-1}\right)}+[g]_{C^{\alpha}\left(\left[t_{0}-4,t_{0}\right]\times S^{N-1}\right)}+\|\varphi\|_{C^{2,\alpha}\left(S^{N-1}\right)}\right],$
we omit the details here. Thus we get the desired estimate by case (i) and case (ii). ∎
@@
and for every $t\in[T,t_{0}]$ and every $k=0,1,2,\ldots,K,$
@@
Proof.
Assume the conclusion is false. Then there exist sequences $\{T_{i}\}$, $\{t_{i}\}$, $\{v_{i}\}$ and $\{g_{i}\}$ with $t_{i}\leq 0$ and $T_{i}-t_{i}\leq-4$, such that
@@
$|\nabla\widetilde{v}_{i}|\leq Ce^{\mu(t_{i}^{*}-T_{i})}\qquad\text{on }(T_{i}-t_{i}^{*},\,T_{i}-t_{i}^{*}+1)\times S^{N-1},$
so
$\left|\tilde{v}_{i}(t,\theta)\right|$ $=\left|\tilde{v}_{i}(0,\theta)-\tilde{v}_{i}\left(T_{i}-t_{i}^{*},\theta\right)\right|$
$\leqslant\sup_{t\in\left(T_{i}-t_{i}^{*},T_{i}-t_{i}^{*}+1\right)}\left|\partial_{t}\tilde{v}_{i}(t,\theta)\right|\left|T_{i}-t_{i}^{*}\right|$
$\leqslant\sup_{(t,\theta)\in\left(T_{i}-t_{i}^{*},T_{i}-t_{i}^{*}+1\right)\times S^{N-1}}\left|\nabla\tilde{V}_{i}(t,\theta)\right|\cdot\left|T_{i}-t_{i}^{*}\right|$
$\leqslant ce^{\mu\left(t_{i}^{*}-T_{i}\right)}\left|T_{i}-t_{i}^{*}\right|$
@@
Also $\widetilde{g}_{i}$ converges uniformly to zero on every compact subset of $(\tau_{-},\tau_{+})\times S^{N-1}$. Therefore, $\widehat{v}\neq 0$, and
$|e^{-\mu t}\widehat{v}(t,\theta)|\leq 1\quad\text{for all}(t,\theta)\in(\tau_{-},\tau_{+})\times S^{N-1}$ (56)
Also we have
$\mathcal{L}\widehat{v}=0\quad on~(\tau_{-},\tau_{+})\times S^{N-1},$
and
$\lim_{t\to\tau}\widehat{v}(t,\theta)=0.$ (57)
Where $\tau=\tau_{-}$ or $\tau_{+}$ if it is finite.
@@
$|e^{-\mu t}\widehat{v}_{k}(t)|\leq C\quad for~all~t\in(\tau_{-},\tau_{+}).$
If $\tau_{+}=+\infty$, then $c_{k}^{1}=0$, thus $\widehat{v}_{k}(t)=c_{k}^{2}e^{\sigma_{2}^{(k)}t}$ which decays exponentially as $t\to+\infty$. If $\tau_{+}$ is finite, then $\lim_{t\to\tau_{+}}\widehat{v}_{k}(t)=0$ by (57).
Similarly, if $\tau_{-}=$-$\infty$, then $c_{k}^{2}=0$, hence $\widehat{v}_{k}(t)=c_{k}^{1}e^{\sigma_{1}^{(k)}t}$ which decays exponentially as $t\to$-$\infty$. If $\tau_{-}$ is finite, then $\lim_{t\to\tau_{-}}\widehat{v}_{k}(t)=0$ by (57).
@@
$\lambda_{k}-(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)>0.$
@@
$\widehat{v}_{k}\equiv 0\qquad\text{for all }k=0,1,\ldots,K\text{ with }\sigma_{1}^{(K)}<\mu.$
Combining the above, we conclude that $\widehat{v}_{k}\equiv 0$ for all $k\geq 0$, therefore $\widehat{v}\equiv 0$ and it contradicts with our earlier conclusion that $\widehat{v}\neq 0$. Lemma 4.3 is proved. ∎
@@
Proof.
For each $k=0,1,\ldots,K$, define
@@
Then
@@
$v_{k}(t)$ $={B_{k}^{1}}^{*}\int_{-\infty}^{t}e^{R(\sigma_{1}^{(k)})(t-s)}g_{k}(s)\sin\bigl(I(\sigma_{1}^{(k)})(t-s)\bigr)\,ds,$
or
$v_{k}(t)=e^{\sigma_{1}^{(k)}t}\int_{-\infty}^{t}\int_{-\infty}^{s}e^{-\sigma_{1}^{(k)}\gamma}g_{k}(\gamma)d\gamma ds,$
where $B_{k}^{1}~and~{B_{k}^{1}}^{*}$ are constants. We now write down the argument for the first and the other are similar.
@@
Combining the above estimates obtains that (58) admits a solution $v_{k}\in C_{\mu}^{2,\alpha}((-\infty,t_{0}])$ satisfying
[…]
where the constant $C$ depends only on $N,\alpha,\mu$ and not on $t_{0}$.
@@
$\|v\|_{C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})}\leq C\sum_{k=0}^{K}\|v_{k}\|_{C_{\mu}^{2,\alpha}((-\infty,t_{0}])}\leq C\sum_{k=0}^{K}\|g_{k}\|_{C_{\mu}^{0,\alpha}((-\infty,t_{0}])}\leq C\|g\|_{C_{\mu}^{0,\alpha}((-\infty,t_{0}]\times S^{N-1})}.$
Thus $v$ is the desired solution. Moreover, it’s clear that such a solution is unique under the additional condition that $v(t,\cdot)\in\operatorname{span}\{Q_{0},Q_{1},\ldots,Q_{K}\}~\text{for every }t\leq t_{0}.$
@@
Proof.
Fix any $T\leq t_{0}-4$. We first show that there exists $v^{\prime}\in C^{2,\alpha}([T,t_{0}]\times S^{N-1})$ such that
$\begin{cases}\mathcal{L}v^{\prime}=g&\text{in }(T,t_{0})\times S^{N-1},\\[2.84526pt] v^{\prime}=0&\text{on }(\{T\}\cup\{t_{0}\})\times S^{N-1}.\end{cases}$ (61)
@@
$\begin{cases}\frac{\partial}{\partial t}\left(e^{{A}t}\frac{\partial v^{\prime}}{\partial t}\right)+e^{{A}t}\Delta_{\theta}v^{\prime}+(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)e^{{A}t}v^{\prime}=e^{{A}t}g&\text{ in }\left(T,t_{0}\right)\times S^{N-1},\\ v^{\prime}=0&\text{ on }\left(\{T\}\cup\{t_{0}\}\right)\times S^{N-1},\end{cases}$
Here ${A}=N-2+\frac{2\alpha+4}{p+1}$. Consider the energy functional
$\mathcal{G}_{T}(v^{\prime})=\int_{T}^{t_{0}}\int_{S^{N-1}}\left[e^{{A}t}(\partial_{t}v^{\prime})^{2}+e^{{A}t}|\nabla_{\theta}v^{\prime}|^{2}-(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)e^{{A}t}{v^{\prime}}^{2}+2e^{{A}t}gv^{\prime}\right]dt\,d\theta.$
@@
$\Gamma=\left\{\phi\in H^{1}(S^{N-1}):\int_{S^{N-1}}\phi(\theta)\,Q_{k}(\theta)\,d\theta=0,\quad k=0,1,2,\ldots,K,\ \sigma_{1}^{(k)}<\mu\right\},$
set $\phi=\sum_{k\geq 0}a_{k}Q_{k}$, then
$\int\phi^{2}=\sum_{k\geq 0}\left|a_{k}\right|^{2},\quad\int\left|\nabla_{\theta}\phi\right|^{2}=\int\phi(-\Delta)\phi=\sum_{k\geq 0}\lambda_{k}\left|a_{k}\right|^{2},$
by the definition of $\Gamma$, we know
$a_{0}=a_{1}=\cdots=a_{K}=0,$
so
$\int\left|\nabla_{\theta}\phi\right|^{2}=\sum_{k\geq K+1}\lambda_{k}\left|a_{k}\right|^{2}\geq\lambda_{K+1}\sum_{k\geq K+1}\left|a_{k}\right|^{2}=\lambda_{K+1}\int\phi^{2}.$
Since $\sigma_{1}^{(K+1)}>\mu>0$, we have
$\lambda_{K+1}-(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)>0,$
it follows that for any $v^{\prime}\in H_{0}^{1}((T,t_{0})\times S^{N-1})$ satisfying $v^{\prime}(t,\cdot)\in\Gamma$ ($t\in(T,t_{0})$), we have
$\mathcal{G}_{T}(v^{\prime})\geq\int_{T}^{t_{0}}\int_{S^{N-1}}e^{{A}t}(\partial_{t}v^{\prime})^{2}+\left(\lambda_{K+1}-(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)\right)e^{{A}v^{\prime}}v^{2}+2e^{{A}v^{\prime}}gv^{\prime}dtd\theta.$
Hence the functional $\mathcal{G}_{T}$ is coercive and weakly lower semicontinuous. Therefore we may find a minimizer $v^{\prime}$ of $\mathcal{G}_{T}$ in the space
$\left\{v^{\prime}\in H_{0}^{1}((T,t_{0})\times S^{N-1}):v^{\prime}(t,\cdot)\in\Gamma\text{ for every }t\in(T,t_{0})\right\}.$
Since $g(t,\cdot)\in\Gamma$ for all $t\in(T,t_{0})$, it follows that $v^{\prime}$ is a solution of (61), and $v^{\prime}(t,\cdot)\in\Gamma$ for all $t\in(T,t_{0})$.
@@
$\sup_{(t,\theta)\in[T,t_{0}]\times S^{N-1}}e^{-\mu t}\,|v^{\prime}(t,\theta)|\leq C\sup_{(t,\theta)\in[T,t_{0}]\times S^{N-1}}e^{-\mu t}\,|g(t,\theta)|,$
@@
For any fixed $T_{0}<t_{0}$, consider the region $t_{0}+T_{0},\,t_{0}]\times S^{N-1}\subset[t_{0}+T_{0}-1,\,t_{0}]\times S^{N-1}$. By the interior and boundary Schauder estimates, together with the fact that $v^{\prime}(t_{0},\theta)=0$, we may extract a subsequence $v^{\prime}$ converges to a $C^{2,\alpha}$ solution $v$ of (52) on $t_{0}+T_{0},\,t_{0}]\times S^{N-1}$ by Arzelà–Ascoli theorem, with $v=0$ on $\{t_{0}\}\times S^{N-1}$ as $T\to$-$\infty$. By a diagonalization process, we conclude that $v^{\prime}$ converges to a $C^{2,\alpha}$ solution $v$ of (52) on $(-\infty,t_{0}]\times S^{N-1}$ , with $v=0$ on $\{t_{0}\}\times S^{N-1}$.
@@
Substituting (62) into (54) (with $\varphi=0$) and then completes the proof. ∎
@@
Proof.
Let $K\geq 0$ be the largest integer such that $\sigma_{1}^{(K)}<\mu$. For $k=0,1,\ldots,K$, define
@@
Let $v_{1}\in C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})$ be the unique solution (as in Lemma 4.4) of
@@
Let $v_{2}\in C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})$ to be the unique solution (as in Lemma 4.6) of
@@
Then $v=v_{1}+v_{2}$ is the desired solution.
Remark.
We denote $\mathcal{L}^{-1}$ the correspondence $g\mapsto v$ in Theorem 4.6. Then
@@
It’s equivalent to prove $H(z)=0$ admits a solution.
@@
$\big|H(\hat{z})\big|+\left|\frac{d}{dt}H(\hat{z})\right|\leq Ce^{\mu t}\quad\forall\,t\in(-\infty,0].$
@@
Proof.
We only prove case (2) since the radial case (1) is similar and easier. For any $\phi\in C_{\mu}^{2,\alpha}((-\infty,t_{0}]\times S^{N-1})$, we have
$H(\widehat{z}+\phi)=H(\widehat{z})+\mathcal{L}\phi-P(\phi),$
@@
Thus $H(\widehat{z}+\phi)=0$ is equivalent to
$\mathcal{L}\phi=-\big[H(\widehat{z})-P(\phi)\big].$ (64)
By Theorem 4.6 and corresponding remark, we may rewrite (63) in the form
$\phi=\mathcal{L}^{-1}\big(-H(\widehat{z})+P(\phi)\big).$
@@
$\mathcal{T}(\phi)=\mathcal{L}^{-1}\big(-H(\widehat{z})+P(\phi)\big).$
@@
$\|H(\widehat{z})\|_{C_{\mu}^{1}((-\infty,t_{0}]\times S^{N-1})}\leq C_{1}.$
@@
(2)Nonradial case. Assume that the corresponding index sets $\mathcal{I}_{\rho}$ and $\mathcal{I}_{\widetilde{\rho}}$ given in (34) and (35) (with multiple number ignored) in the proof of Theorem 2.2 respectively. $\mu\notin\mathcal{I}_{\rho}\cup\mathcal{I}_{\widetilde{\rho}}~\text{and}~\mu>\rho_{j+4}>\frac{2p-\alpha}{p+1}$ in cases (31)(32); $\mu>\rho_{1}$in cases (30)(33). Let $\varphi$ be a solution of the equation $\mathcal{L}(\varphi)=0$ on $\mathbb{R}\times\mathbb{S}^{N-1}$, and suppose that $\varphi(t,\theta)\to 0~~\text{as}~t\to-\infty~\text{uniformly}~\text{for}~\theta\in{S^{N-1}}$. Then there exists $t_{0}<0$ with $|t_{0}|$ sufficiently large, such that there exists a smooth function $\widetilde{\varphi}$ on $(-\infty,t_{0}]\times\mathbb{S}^{N-1}$ satisfying
@@
Proof.
We prove case (2); the proof of case (1) is similar and easier.